# Cosmography with next-generation gravitational wave detectors

### v1

@@ Abstract
Advancements in cosmology through next-generation ground-based gravitational wave observatories will bring in a paradigm shift. We explore the pivotal role that gravitational-wave standard sirens will play in inferring cosmological parameters with next-generation observatories, not only achieving exquisite precision but also opening up unprecedented redshifts. We examine the merits and the systematic biases involved in gravitational-wave standard sirens utilizing binary black holes, binary neutron stars, and neutron star-black hole mergers. Further, we estimate the precision of bright sirens, golden dark sirens, and spectral sirens for these binary coalescences and compare the abilities of various next-generation observatories (${}^{\sharp}$, Cosmic Explorer, Einstein Telescope, and their possible networks). When combining different sirens, we find sub-percent precision over more than 10 billion years of cosmic evolution for the Hubble expansion rate $H(z)$. This work presents a broad view of opportunities to precisely measure the cosmic expansion rate, decipher the elusive dark energy and dark matter, and potentially discover new physics in the uncharted Universe with next-generation gravitational-wave detectors.
  • February 5, 2024
@@ 1 Introduction
In the next decades, gravitational-wave (GW) observations are expected to uncover many mysteries in the Universe. The planned next-generation ground-based and space-based GW observatories will bring a wealth of science opportunities with their reaches to the early Universe [1, 2, 3, 4, 5, 6, 7]. The current generation of ground-based GW observatories, LIGO, Virgo, and KAGRA, with their successive upgrades (e.g., A+ and ${}^{\sharp}$) will be able to detect stellar-mass binary black hole (BBH) mergers up to redshift $z\sim 10$ [8]. Next-generation (XG) ground-based GW detectors, Cosmic Explorer (CE) and Einstein Telescope (ET), aim at observing binary neutron star (BNS) and neutron star-black hole (NSBH) mergers up to $z\sim 10$ as well as all stellar-mass BBHs in the Universe ($z\sim 100$). These observations of mergers across cosmic time will make XG detectors new powerful cosmological probes, mapping the Universe from the present dark energy domination to the past dark matter era, cf. figure 1.
In this era of precision cosmology with many well-established cosmological measurements, GW observations have several unique roles to play. First, GW observations will independently measure the cosmological parameters, with the potential to resolve tensions among existing electromagnetic (EM) measurements [9]. Second, the detections across different redshifts allow for the mapping of the Universe’s expansion history with a single probe over more than 10 billion years of cosmic evolution. Third, with the Universe observed in GWs for the first time, XG detectors will verify the consistency between the EM and GW cosmological measurements and reveal (if any) underlying new physics. These observations will test gravity in new regimes, reaching unprecedented strong fields and cosmic scales.
@@
Assuming general relativity, the distances to a given binary can be determined from the GW signals released from the binary; meanwhile, multiple techniques are available to estimate the redshift of the binary. With the distance and redshift estimate, GW sources can be used to measure the Universe expansion rate $H(z)$ and construct the GW Hubble diagram. In analogy to the use of supernovae Type Ia as standard candles for cosmology [10, 11], this is known as the ‘standard siren’ method [12, 13]. The observations of BNS GW170817 and its EM counterparts enabled the first standard siren measurement [14], opening a new era in multi-messenger astronomy (MMA). After the method was proposed in 1986 [12], new ideas on redshift acquirement, developments in data analysis, applications to GW data, and investigations of systematic uncertainties promised an exciting path forward for standard siren measurements. In this article, we explore further the prospects of the standard siren cosmography in the XG era.
@@ 2 Standard siren cosmology
Gravitational waves are natural messengers to probe the cosmic history. Unlike most other transients in the Universe, their signals can be well-understood from first principles, as general relativity has clear predictions for the waveform of a compact binary coalescence. Moreover, GW detectors are directly sensitive to the waveform of the signal (as opposed to the intensity) allowing us to observe signals further away, since the GW amplitude scales with the inverse of the luminosity distance (and not the inverse square of the distance). Since GWs are coherently detected, the absolute value of the luminosity distance is directly measured—giving GW sources the name of standard sirens. The GW strain, $h_{\mathrm{gw}}$, thus contains information about the expansion rate through the luminosity distance
$h_{\mathrm{gw}}\propto 1/d_{L}\,,\qquad d_{L}=(1+z)\int_{0}^{z}\frac{c\mathrm{d}z^{\prime}}{H(z^{\prime})}\,,$ (1)
where we have assumed a flat background cosmology. Under that assumption and focusing on a Universe dominated by dark energy and dark matter, the Hubble parameter measuring the expansion rate can be written as
@@
where $H_{0}$ is the Hubble constant and denotes the local expansion rate $H_{0}=H(z=0)$, and $\Omega_{\Lambda}$ and $\Omega_{m}$ are the fractional energy density of dark energy and dark matter that, in a flat Universe, satisfy $\Omega_{\Lambda}+\,\Omega_{m}=1$.
@@
What GWs do not provide, however, is a direct account of the precise time when the signal merged. Times, or redshifts, together with distances allow the measurement of the Universe’s expansion rate at different epochs. Therefore, the entire game of GW cosmology is to obtain redshift information. In the following we review different methods to achieve this goal.
@@ 2.1 Bright sirens
Kilonova AT 2017gfo and GRB 170817A [24, 25, 26, 27, 28, 29, 30, 31, 32, 33] were the first EM counterparts found to be associated with a GW source, GW170817 [34]. With the precise localization, the host galaxy NGC 4993 was identified and the redshift could be determined ($z\sim 0.009$). This event led to a Hubble constant measurement of $H_{0}=70^{+12.0}_{-8.0}$km/s/Mpc [14]. Over last few years, a variety of other types of EM emissions were proposed theoretically [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] or suggested observationally [56, 57, 58, 59] to be associated with BNS, NSBH, and BBH mergers. If confirmed, the applications of the bright siren approach will be further widened [60, 61, 62].
If we denote the data from event $i$ as $\vec{\mathcal{D}}_{i}=\{\vec{\mathcal{D}}_{{\rm GW},i},\vec{\mathcal{D}}_{{\rm EM},i}\}$, where $\vec{\mathcal{D}}_{{\rm GW},i}$ and $\vec{\mathcal{D}}_{{\rm GW},i}$ represent the GW and EM data respectively, the likelihood of the bright siren inference can be written as:
[…]
where $\vec{\Theta}$ represents the physical parameters of the binary, such as the luminosity distance, redshift, and sky location. $\vec{\mathcal{D}_{\rm th}}$ denotes the detection threshold for the data. With Advanced LIGO-Virgo-KAGRA (LVK) sensitivities, the Hubble constant can be measured to $\sim 2\%$ when combining $\sim 50$ BNS bright sirens [63, 64, 65, 66]. This precision could shed light on the tension between different Hubble constant measurements [9].
@@ 2.2 Dark sirens
The statistical dark siren method exploits galaxy catalogs to identify the galaxies that lie within the localization volume associated with the GW signal (note that this localization volume is obtained purely from the GW detection [90]) and obtain their redshifts. Together with the probability distribution of $d_{L}$ inferred from the GW signal, the redshifts for each of these potential hosts give corresponding probability distributions on cosmological parameters. The uncertainty in the knowledge of the true host galaxy is accounted for by statistically averaging over the redshifts from all potential host galaxies, which results in constraints on cosmological parameters using a single GW event. Such an analysis can be performed for multiple GW events, which will result in the inference of cosmological parameters with greater precision [64, 91, 92].
@@
GW170814 [93] was the first BBH merger to which this technique was applied. Using the Dark Energy Survey data [94] for redshift information, $H_{0}$ was found to be $75^{+40}_{-32}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ [95]. When applied to the eight best-localized events in GWTC-3, stronger constraints are obtained on $H_{0}$, measured to be $79.8^{+19.1}_{-12.8}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ [96]. The statistical siren approach can be applied to bright siren events as well. However, in general, this method is expected to provide weaker constraints than the bright siren technique due to the uncertainty in host galaxy identification. Ref. [97] applies the statistical siren strategy to GW170817 to obtain $H_{0}=76^{+48}_{-23}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$. But, as mentioned in section 2.1, due to unique host galaxy identification made possible by EM counterpart detection, the bright siren method provides much stronger bounds on $H_{0}$. The constraints obtained from the statistical dark siren and the bright siren approaches were combined for events in GWTC-3 to estimate $H_{0}$ with $\sim 15\%$ uncertainty [98, 96]. Note that this measurement is dominated by the bright siren constraints from GW170817.
@@
Among these dark siren mergers, there may be some that are so well localized in the sky that only one potential host galaxy is found in that localization volume [99, 100]. This would allow unique identification of the host galaxy and lead to a very precise estimation of cosmological parameters. Such mergers are referred to as golden dark sirens. Asymmetric (i.e., the masses of the compact objects are appreciably different) and face-off (i.e., the inclination angle $(\iota)$ is not zero) systems are touted to be ideal candidates for such sirens [101, 102]. This is due to the excitation of higher-order modes in the GW waveform for asymmetric and face-off systems [103], leading to the breaking of the degeneracy between $d_{L}-\iota$, which leads to the precise estimation of parameters [104, 105, 106, 107]. Similar expectations hold for precessing systems. Because golden dark sirens are typically the loudest events, their measurement of the luminosity distance will be extremely precise, outperforming those of bright sirens.
@@ 2.3 Spectral sirens
For the same source population, different expansion rates will predict different distributions of detector frame masses. For instance, as schematically shown in figure 2, a larger value of the local expansion rate $H_{0}$ will imply a larger redshift in order to end with the same luminosity distance, making the detector frame distribution shift faster to higher values. Therefore, one can constrain cosmological parameters with the population analysis of spectral sirens. Since the cosmological redshift is a common effect to all mergers at equal distance, the better the mass spectrum can be constrained over a larger range of distances, the more information can be extracted about cosmology.
@@
The spectral siren method is in its own nature data driven. It does not matter how the binaries are formed or what they are made of, it only matters if a global shift of the mass spectrum can be identified over a wide range of distances. In this sense all mergers can be considered simultaneously and cosmology is constrained with the full mass distribution of compact binaries [131]. This is advantageous because different sectors of the mass spectrum are expected to evolve differently with redshift. For example, the masses of neutron stars and low mass black holes are thought to be much less sensitive to the environment that those of heavier black holes whose possible maximum mass depends significantly on metalicity. Therefore, by studying them together one can break the possible degeneracies between the shift induced by the cosmological redshift and the change in the shape of the mass spectrum due to the time and environment dependence of the source population [131]. If unaccounted for, however, the evolution of the mass spectrum can bias the inference of cosmological parameters [131, 132, 133] as well as the mismodeling of the true distribution with oversimplified parametric models [133]. This is why the non-parametric reconstruction of the mass spectrum stand out as a promising avenue for spectral siren cosmology [134].
It is important to note that our sensitivity to different sectors of the compact binary mass spectrum will evolve in parallel to the upgrades of the GW detector network. While currently the inference of cosmology is dominated by the excess of BBHs at $\sim 35M_{\odot}$ and the lack of them at $>45M_{\odot}$, spectral siren cosmology with XG detectors will be dominated by BNSs and low-mass BBHs, as their intrinsic merger rate is much larger [131]. Different populations will uncover the expansion rate at different redshifts, as their intrinsic merger rate history and detectability highly depends on their masses. Interestingly, with detectors at A+ sensitivity and more so with XG, we could also detect BBHs on the far-side side of the PISN gap (if they exist), thereby resolving the upper edge of the PISN gap and providing an extra anchor to probe cosmology [135].

### v2

@@ Abstract
Advancements in cosmology through next-generation ground-based gravitational wave observatories will bring in a paradigm shift. We explore the pivotal role that gravitational-wave standard sirens will play in inferring cosmological parameters with next-generation observatories, not only achieving exquisite precision but also opening up unprecedented redshifts. We examine the merits and the systematic biases involved in gravitational-wave standard sirens utilizing binary black holes, binary neutron stars, and neutron star-black hole mergers. Further, we estimate the precision of bright sirens, golden dark sirens, and spectral sirens for these binary coalescences and compare the abilities of various next-generation observatories (A, Cosmic Explorer, Einstein Telescope, and their possible networks). When combining different sirens, we find sub-percent precision over more than 10 billion years of cosmic evolution for the Hubble expansion rate $H(z)$. This work presents a broad view of opportunities to precisely measure the cosmic expansion rate, decipher the elusive dark energy and dark matter, and potentially discover new physics in the uncharted Universe with next-generation gravitational-wave detectors.
  • July 16, 2024
@@ 1 Introduction
In the next decades, gravitational-wave (GW) observations are expected to uncover many mysteries in the Universe. The planned next-generation ground-based and space-based GW observatories will bring a wealth of science opportunities with their reaches to the early Universe. The current generation of ground-based GW observatories, LIGO [1], Virgo [2], and KAGRA [3], with their successive upgrades (e.g., A+ [4, 5] and A[6]) will be able to detect stellar-mass binary black hole (BBH) mergers up to redshift $z\sim 10$ [7]. Next-generation (XG) ground-based GW detectors, Cosmic Explorer (CE) [8] and Einstein Telescope (ET) [9], aim at observing binary neutron star (BNS) and neutron star-black hole (NSBH) mergers up to $z\sim 10$ as well as all stellar-mass BBHs in the Universe ($z\sim 100$). [^1] NEMO is another possible future ground-based detector with sensitivity similar to A+ but ranging all the way to $\sim 4$kHz [10]. Therefore, NEMO is optimized to detect the (post-)merger of BNS. We do not discuss NEMO in this work. These observations of mergers across cosmic time will make XG detectors new powerful cosmological probes, mapping the Universe from the present dark energy domination to the past dark matter era, cf. figure 1. Future GW detectors in space will also bring relevant cosmological constraints [11, 12, 13, 14], but we focus on ground-based detectors only here.
In this era of precision cosmology with many well-established cosmological measurements, GW observations have several unique roles to play. First, GW observations will independently measure the cosmological parameters, with the potential to resolve tensions among existing electromagnetic (EM) measurements [15]. Second, the detections across different redshifts allow for the mapping of the Universe’s expansion history with a single probe over more than 10 billion years of cosmic evolution. Third, XG detectors will verify the consistency between the EM and GW cosmological measurements and reveal (if any) underlying new physics. These observations will test gravity in new regimes, reaching unprecedented strong fields and cosmic scales.
@@
Assuming general relativity, the distances to a given binary can be determined from the GW signals released from the binary; meanwhile, multiple techniques are available to estimate the redshift of the binary. With the distance and redshift estimate, GW sources can be used to measure the Universe expansion rate $H(z)$ and construct the GW Hubble diagram. In analogy to the use of supernovae Type Ia as standard candles for cosmology [16, 17], this is known as the ‘standard siren’ method [18, 19]. The observations of the BNS GW170817 and its EM counterparts enabled the first standard siren measurement [20], opening a new era in multi-messenger astronomy (MMA). After the method was proposed in 1986 [18], new ideas on redshift acquirement, developments in data analysis, applications to GW data, and investigations of systematic uncertainties promised an exciting path forward for standard siren measurements. In this article, we explore further the prospects of the standard siren cosmography in the XG era.
@@ 2 Standard siren cosmology
Gravitational waves are natural messengers to probe the cosmic history. Unlike most other transients in the Universe, their signals can be well-understood from first principles, as general relativity has clear predictions for the waveform of a compact binary coalescence. For example, at leading order, the frequency evolution is determined by a specific combination of the detected component masses of the binary known as the chirp mass $\mathcal{M}_{c}=(m_{1}m_{2})^{3/5}/(m_{1}+m_{2})^{1/5}$: $\dot{f}_{\mathrm{gw}}\propto\mathcal{M}_{c}^{5/3}f_{\mathrm{gw}}^{11/3}$ [25]. Moreover, GW detectors are directly sensitive to the waveform of the signal (as opposed to the intensity) allowing us to observe signals further away, since the GW amplitude scales with the inverse of the luminosity distance (and not the inverse square of the distance). Since GWs are coherently detected, the absolute value of the luminosity distance is directly measured—giving GW sources the name of standard sirens. The GW strain, $h_{\mathrm{gw}}$, thus contains information about the expansion rate through the luminosity distance
$h_{\mathrm{gw}}\propto{\mathcal{M}_{c}^{5/3}}/d_{L}\,,\qquad d_{L}=(1+z)\int_{0}^{z}\frac{c\mathrm{d}z^{\prime}}{H(z^{\prime})}\,,$ (1)
where we have assumed a flat background cosmology. Under that assumption and focusing on a Universe dominated by dark energy and dark matter (radiation is only relevant in the early Universe), the Hubble parameter measuring the expansion rate can be written as
@@
where $H_{0}$ is the Hubble constant and denotes the local expansion rate $H_{0}=H(z=0)$, and $\Omega_{\Lambda}$ and $\Omega_{m}$ are the fractional energy density of dark energy and dark matter that, in a flat Universe, satisfy $\Omega_{\Lambda}+\,\Omega_{m}=1$. The latest cosmological observations from the Planck satellite result in $H_{0}=67.66\pm 0.42$km/s/Mpc, $\Omega_{\Lambda}=0.6889\pm 0.0056$ and $\Omega_{m}=0.3111\pm 0.0056$ [26]. Alternative cosmological models are also possible, potentially modifying both the background cosmology $H(z)$ and the scaling of the strain with the luminosity distance. We will discuss those in section 4.5.
@@
What GWs do not provide, however, is a direct account of the precise epoch when the binary merged. Cosmic times, or redshifts, together with distances allow the measurement of the Universe’s expansion rate at different epochs. Therefore, the entire game of GW cosmology is to obtain redshift information. In the following we review different methods to achieve this goal.
@@ 2.1 Bright sirens
Kilonova AT 2017gfo and GRB 170817A [32, 33, 34, 35, 36, 37, 38, 39, 40, 41] were the first EM counterparts found to be associated with a GW source, GW170817 [42]. With the precise localization, the host galaxy NGC 4993 was identified and the redshift could be determined ($z\sim 0.009$). This event led to a Hubble constant measurement of $H_{0}=70^{+12.0}_{-8.0}$km/s/Mpc [20]. With Advanced LIGO-Virgo-KAGRA (LVK) sensitivities, the Hubble constant can be measured to $\sim 2\%$ when combining $\sim 50$ BNS bright sirens [43, 44, 45, 46]. This precision could shed light on the tension between different Hubble constant measurements [15]. Over last few years, a variety of other types of EM emissions were proposed theoretically [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] or suggested observationally [68, 69, 70, 71] to be associated with BNS, NSBH, and BBH mergers, although they could not be confidently verified. If confirmed, the applications of the bright siren approach will be further widened [72, 73, 74].
@@ 2.2 Dark sirens
The statistical dark siren method exploits galaxy catalogs to identify the galaxies that lie within the localization volume associated with the GW signal (note that this localization volume is obtained purely from the GW detection [98]) and obtain their redshifts. Together with the probability distribution of $d_{L}$ inferred from the GW signal, the redshifts for each of these potential hosts give corresponding probability distributions on cosmological parameters. The uncertainty in the knowledge of the true host galaxy is accounted for by statistically averaging over the redshifts from all potential host galaxies, which results in constraints on cosmological parameters using a single GW event. Such an analysis can be performed for multiple GW events, which will result in the inference of cosmological parameters with greater precision [44, 99, 100, 101].
@@
GW170814 [102] was the first BBH merger to which this technique was applied. It was localized to a volume of $2\times 10^{6}\,\mathrm{Mpc}^{3}$ [103] which, according to the Dark Energy Survey (DES) data, contained $\sim 77,000$ galaxies [104]. In comparison, the localization volume of the best localized BBH, GW190814 [105], was $\sim 100$ times smaller and contained $\sim 2000$ galaxies [106]. Using GW170814 and the DES data [107] for redshift information, $H_{0}$ was found to be $75^{+40}_{-32}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ [104]. When applied to the eight best-localized events in GWTC-3, stronger constraints are obtained on $H_{0}$, measured to be $79.8^{+19.1}_{-12.8}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ [103]. The statistical siren approach can be applied to bright siren events as well. However, in general, this method is expected to provide weaker constraints than the bright siren technique due to the uncertainty in host galaxy identification. Ref. [108] applies the statistical siren strategy to GW170817 to obtain $H_{0}=76^{+48}_{-23}\,\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$. But, as mentioned in section 2.1, due to unique host galaxy identification made possible by EM counterpart detection, the bright siren method provides much stronger bounds on $H_{0}$. The constraints obtained from the statistical dark siren and the bright siren approaches were combined for events in GWTC-3 to estimate $H_{0}$ with $\sim 15\%$ uncertainty [109, 103]. Note that this measurement is dominated by the bright siren constraints from GW170817.
@@
The advent of XG observatories will see an improvement in the measurement precision of both luminosity distance and sky-area [110, 111, 112, 113, 7]. This will result in smaller localization volumes containing smaller number of potential hosts compared to currently detected events [114]. Among these dark siren mergers, there may be some that are so well localized in the sky that only one potential host galaxy is found in that localization volume [115, 116]. This would allow unique identification of the host galaxy and lead to a very precise estimation of cosmological parameters. Such mergers are referred to as golden dark sirens. Asymmetric (i.e., the masses of the compact objects are appreciably different) and face-off (i.e., the inclination angle $(\iota)$ is not zero) systems are touted to be ideal candidates for such sirens [117, 118]. This is due to the excitation of higher-order modes in the GW waveform for asymmetric and face-off systems [119, 120, 121], leading to the breaking of the degeneracy between $d_{L}-\iota$, which leads to the precise estimation of parameters [122, 123, 124, 125, 126, 127]. Similar expectations hold for precessing systems in which the spins are not (anti-)aligned and there could be precession of the orbital plane [128, 127, 129]. Because golden dark sirens are typically the loudest events, their measurement of the luminosity distance will be extremely precise, outperforming those of bright sirens.
@@ 2.3 Spectral sirens
For the same source population, different expansion rates will predict different distributions of detector frame masses. For instance, as schematically shown in figure 2, a larger value of the local expansion rate $H_{0}$ will imply a larger redshift in order to end with the same luminosity distance, making the detector frame distribution shift faster to higher values. Therefore, one can constrain cosmological parameters with the population analysis of spectral sirens. Since the cosmological redshift is a common effect to all mergers at equal distance, the better the mass spectrum can be constrained over a larger range of distances, the more information can be extracted about cosmology. One should be cautious, however, because the mass spectrum is expected to evolve with redshift and this could bias the cosmological inference if not properly accounted for, as we discuss later.
@@
The spectral siren method is in its own nature data driven: the cosmology is jointly inferred with the population properties which are directly obtained from the data. It does not matter how the binaries are astrophysically formed or what they are made of, it only matters if a global shift of the mass spectrum can be identified over a wide range of distances and distinguished from local changes of its shape. In this sense all mergers can be considered simultaneously, from light neutron stars to heavy black holes, and cosmology is constrained with the full mass distribution of compact binaries [152]. This is advantageous because different sectors of the mass spectrum are expected to evolve differently with redshift. For example, the masses of neutron stars and low mass black holes are thought to be much less sensitive to the environment than those of heavier black holes whose possible maximum mass depends significantly on metalicity. Therefore, by studying the whole mass spectrum together, with its many individual features, one can break the possible degeneracies between the shift induced by the cosmological redshift and the change in the shape of the mass spectrum due to the time and environment dependence of the source population [152]. If unaccounted for, however, the evolution of the mass spectrum can bias the inference of cosmological parameters [152, 153, 154], a potential problem especially relevant for XG detectors observing far away binaries. Mismodeling of the true distribution with oversimplified parametric models [154] or wrong astrophysical assumptions that do not fit well the data can also be problematic. This is why the non-parametric reconstruction of the mass spectrum stand out as a promising avenue for spectral siren cosmology with future observations, bypassing the need to choose a specific mass function and just relying on shape extracted from the (large) data sets [155].
Finally, it is important to note that our sensitivity to different sectors of the compact binary mass spectrum will evolve in parallel to the upgrades of the GW detector network. While currently the inference of cosmology is dominated by the excess of BBHs at $\sim 35M_{\odot}$ and the lack of them at $>45M_{\odot}$, spectral siren cosmology with XG detectors will be dominated by BNSs and low-mass BBHs, as their intrinsic merger rate is much larger [152]. Different populations will uncover the expansion rate at different redshifts, as their intrinsic merger rate history and detectability highly depends on their masses. Interestingly, with detectors at A+ sensitivity and more so with XG, we could also detect BBHs on the far-side side of the PISN gap (if they exist), thereby resolving the upper edge of the PISN gap and providing an extra anchor to probe cosmology [156].